# Guiding diffusion models to reconstruct flow fields from sparse data

### v1

@@ III.1 Test cases
For the 2D Kolmogorov flow, we follow the configuration of the previous work Shu, Li, and Farimani (2023). The system is governed by the following Navier-Stokes equation in its vorticity form:
@@
where $\textbf{x}\in(0,2\pi)^{2}$, $\omega$ is the vorticity, $\mathbf{u}$ the velocity field and $f(\mathbf{x})=-4\cos(4x_{2})-0.1\omega(\mathbf{x},t)$ is the external force. The Reynolds number $Re$ is set to 1000. The simulation data contains 4240 samples of size $256\times 256$, from which 3000 are randomly selected for training and the remaining 1240 to test the models. To evaluate the physical consistency of the generated samples, we compute the residual of equation (17), where the time derivative is approximated with three generated consecutive timesteps as $\partial_{t}~\omega~\approx~(\omega(\mathbf{x},t-\Delta t)+\omega(\mathbf{x},t+\Delta t))/(2\Delta t)$, and the spatial derivates are calculated with the Fourier pseudo-spectral method.
For the 3D turbulent scenario, we choose a forced isotropic turbulence simulation from the Johns Hopkins Turbulence Database (JHTDB) Databases (2019). It consists of a DNS simulation with a pseudo-spectral method using 10243 nodes in a periodic domain of size $(0,2\pi)^{3}$ during $T=10$ s using 5028 timesteps. Energy is injected to keep the total energy constant in modes such that their wave number magnitude is less than or equal to 2. We use a smaller derived dataset, taking 500 snapshots at a smaller resolution of $128^{3}$ using spectral downsampling. Finally, we shuffle all snapshots and use 400 and 100 samples for training and testing, respectively. The divergence of the flow field $\nabla\cdot\mathbf{u}$ is used as the metric for physics consistency. Figures 4(a) and 4(b) show samples of our training data.
@@ III.2 Physcial-consistency
We first target an unconditioned generation task to examine the ability of different methods to generate physically constrained samples. We consider standard diffusion models, Physics-Informed Diffusion Models Bastek, Sun, and Kochmann (2024) with dynamic rescale factor ($\text{PIDM}_{\text{dyn}}$), ConFIG Liu, Chu, and Thuerey (2024), and our $\text{ConFIG}_{\text{u}}$. Some comments about training hyperparameters and the network’s architecture can be found in Appendix B.
To calculate the physical residual, the time and spatial derivatives are evaluated in a discrete manner. Ideally, the training data’s physical residual $\|\mathcal{R}(\omega^{*})\|_{2}$ should be 0. Still, we usually encounter a non-zero physical residual due to the discretisation errors. As no method can be expected to completely resolve this discretization error, it obscures the capabilities of the methods to improve the diffusion training. Hence, we employ a metric that compares the residual of the generated samples $\omega^{\prime}$ to the average of the test dataset samples:
@@
$\text{PIDM}_{\text{dyn}}$ Bastek, Sun, and Kochmann (2024) 0.99 $\pm$ 0.21 0.316 $\pm$ 0.005
ConFIG Liu, Chu, and Thuerey (2024) 1.24 $\pm$ 0.32 0.340 $\pm$ 0.009
@@
$\text{PIDM}_{\text{dyn}}$ Bastek, Sun, and Kochmann (2024) 1.08 $\pm$ 0.07 0.696 $\pm$ 0.042
ConFIG Liu, Chu, and Thuerey (2024) 1.01 $\pm$ 0.05 0.667 $\pm$ 0.067
@@ III.3 Reconstruction tasks
Next, we test the proposed masked diffusion method for reconstructing high-fidelity flow fields from sparse data points. Following the configuration from the previous study by Shu et al. Shu, Li, and Farimani (2023), we target reconstruction tasks with sparse observations of 5% and 1.5625% of the high fidelity data.
@@
The first property is typically measured using the Root Mean Square Error (RMSE) between the predicted flow field and the ground truth. To satisfy physical coherence, the reconstructed field should follow the physical residual of the data, which can be evaluated using the relative residual loss $\boldsymbol{L_{\text{{res}}}}$ introduced in Section III.2.
Evaluating the structure of a fluid flow improves our understanding of how closely the predicted sample matches the target high-fidelity field. As no ground truth is available in this probabilistic setting, we use a learned, perceptual similarity metric (LSiM) Kohl, Um, and Thuerey (2020) to measure the similarity between our reconstruction and the ground truth. This metric allows us to assess whether the structural characteristics of the generated sample align with those of the target field.
@@
For the 2D scenario, we compute this error using the enstrophy spectrum instead of the energy spectrum as used in other works Chen et al. (2024), since enstrophy cascades to small scales in 2D and therefore provides a more sensitive and physically meaningful diagnostic of whether the simulation captures the correct multiscale turbulent dynamics.
@@
Besides comparing with the method of Shu et al. Shu, Li, and Farimani (2023), we also compare our reconstruction result with the supervised baseline where the network directly predicts the high-fidelity data from an interpolated sparse input. For a fair comparison, we use the same backbone architectures as the ones for the diffusion models. Moreover, we need to highlight that the baseline and other diffusion models for sparse data reconstruction require specific training for each new scenario (different % of data, for instance) to maximize their performance. In contrast, the masked diffusion method is general and only needs to be trained once, as an unconditional generator, using only the high-resolution training samples. Table 2 shows the results of the 2D 5% reconstruction task, where we can observe how our masked diffusion method outperforms previous methods on the point-wise, residual, structure similarity, and energy spectrum errors. Similar results can be observed when evaluating the 1.5625% task, which can be seen in Table 3. Our method consistently obtains the lowest errors in both tasks.
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.79 $\pm$ 0.09 0.16 $\pm$ 0.07 0.29 $\pm$ 0.02 0.46 $\pm$ 0.12
Shu et al. Shu, Li, and Farimani (2023) 0.18 $\pm$ 0.03 0.84 $\pm$ 0.07 0.10 $\pm$ 0.02 0.54 $\pm$ 0.08
Masked diffusion (ours) 0.13 $\pm$ 0.02 0.02 $\pm$ 0.03 0.05 $\pm$ 0.01 0.26 $\pm$ 0.05
@@
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 1.06 $\pm$ 0.16 0.21 $\pm$ 0.14 0.42 $\pm$ 0.06 0.68 $\pm$ 0.23
Shu et al. Shu, Li, and Farimani (2023) 0.37 $\pm$ 0.06 0.22 $\pm$ 0.11 0.17 $\pm$ 0.03 0.52 $\pm$ 0.12
Masked diffusion (ours) 0.31 $\pm$ 0.06 0.15 $\pm$ 0.06 0.11 $\pm$ 0.02 0.33 $\pm$ 0.07
@@
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.13 $\pm$ 0.01 0.32 $\pm$ 0.02 0.05 $\pm$ 0.01 0.64 $\pm$ 0.02
Shu et al. Shu, Li, and Farimani (2023) 0.45 $\pm$ 0.03 0.80 $\pm$ 0.06 0.64 $\pm$ 0.15 0.38 $\pm$ 0.07
Masked diffusion (ours) 0.17 $\pm$ 0.01 0.61 $\pm$ 0.03 0.06 $\pm$ 0.01 0.09 $\pm$ 0.01
@@
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.16 $\pm$ 0.01 0.36 $\pm$ 0.01 0.08 $\pm$ 0.01 0.71 $\pm$ 0.02
Shu et al. Shu, Li, and Farimani (2023) 0.44 $\pm$ 0.02 0.79 $\pm$ 0.05 0.62 $\pm$ 0.11 0.36 $\pm$ 0.04
Masked diffusion (ours) 0.20 $\pm$ 0.01 0.66 $\pm$ 0.03 0.09 $\pm$ 0.01 0.11 $\pm$ 0.01
@@
In Tables 4 and 5 we show the results for the 3D case of the same two reconstruction tasks. Our masked diffusion method outperforms Shu et al.’s method across all evaluation metrics. However, the supervised baseline achieves lower errors in RMSE, physical residual, and LSiM. These three metrics are primarily influenced by the accuracy of the low-frequency components, which the supervised baseline captures more effectively. As shown in Figure 7, only the diffusion methods are capable of accurately reconstructing the high-frequency components. In contrast, the supervised baseline produces noticeably blurrier results despite achieving a lower RMSE. The spectrum in Figure 7 shows that the high wavenumber content of the solutions inferred by the supervised baseline clearly lacks energy.
@@
Figures 8-11 show samples of the reconstruction in 2D and 3D cases. It can be seen how masked diffusion better approximates the complicated structures of the turbulent flow fields, producing sharper images and yielding a closer approximation of the flow structures and statistics.
Figure 8: Example 2D Kolmogorov reconstruction 5%
Figure 9: Example 2D Kolmogorov reconstruction 1.5625%
@@
As is evident from the spectrum errors in Tables 4 and 5, only the two considered diffusion models are capable of capturing small-scale features. E.g., our method features more than $6.4\times$ lower error than the supervised baseline. However, the two diffusion approaches behave very differently when considering different amounts of available data. In Figure 12, we plot the RMSE over the 3D reconstruction of 50 samples with Shu et al.’s method and our masked diffusion method together with cubic interpolation. We observe that Shu et al.’s method remains mostly constant, with a consistently higher RMSE than our method. We see that if the percentage of available data is too low, our method achieves even lower RMSE errors than cubic interpolation, in addition to producing high-frequency content.
@@
We also study the distributional accuracy of the methods. For the 2D experiment, we measure the distribution of the vorticity magnitude of the reconstructed samples, as shown in Figure 13. Our method closely approximates the reference’s vorticity distribution on the 5% task, while Shu et al.’s method slightly underestimates the magnitude of the generated samples Shu, Li, and Farimani (2023). When evaluating the 1.5625% task, our method also underestimates the vorticity, but gives a better distribution than Shu et al.’s method.
@@
In Figure 14, we clearly observe that our masked diffusion is the closest to the ground truth, while the supervised baseline overestimates the zero component. Figures 15 and 16 show visualizations of the Q-criterion, highlighting how our proposed method is the only one that captures most of the smaller vortices in the fluid.
When discussing efficiency, our method obtains these results using fewer reverse steps than Shu et al.Shu, Li, and Farimani (2023), which makes it a more computationally efficient model. For instance, for the more demanding 3D experiments in a NVIDIA L40S GPU, as shown in Appendix C, 100 reverse steps, i.e. NN evaluations, are enough to achieve maximum performance, compared to the T=1000 during training, or the 300 reverse steps used with Shu et al.’s method. This also results in shorter running times, where the supervised baseline needs around 0.04 s for a single sample, Shu et al.’s method 25 s, and our method 7.7 s. The only inconvenience of our method would be higher memory demands. During masked diffusion, we need to store the mask and the interpolated sparse data, resulting in 5.2 GB, compared with the 4.0 GB of Shu et al.’s method, and 1.4 GB of the single-step supervised baseline.
Overall, our proposed training and sampling techniques demonstrate a clearly improved performance compared to existing approaches. These results highlight the ability of our masked diffusion framework to capture fine-grained details and preserve the structural integrity of the reconstructed fluid fields.
@@
Figure 15: Q-citerion in 3D for the 5% sparse data reconstruction. 3D plots showing Q isosurfaces with vorticity coloring.
Figure 16: Q-citerion in 3D for the 1.5625% sparse data reconstruction. 3D plots showing Q isosurfaces with vorticity coloring.

### v2

@@ III.1 Test cases
For the 2D Kolmogorov flow, we follow the configuration of the previous work Shu et al. (2023). The system is governed by the following Navier-Stokes equation in its vorticity form:
@@
where $\textbf{x}\in(0,2\pi)^{2}$, $\omega$ is the vorticity, $\mathbf{u}$ the velocity field and $f(\mathbf{x})=-4\cos(4x_{2})-0.1\omega(\mathbf{x},t)$ is the external force. The Reynolds number $Re$ is set to 1000. The simulation data contains 4240 samples of size $256\times 256$, from which 3000 are randomly selected for training and the remaining 1240 to test the models. To evaluate the physical consistency of the generated samples, we compute the residual of equation (17), where the time derivative is approximated with three generated consecutive timesteps as $\partial_{t}~\omega~\approx~(\omega(\mathbf{x},t-\Delta t)+\omega(\mathbf{x},t+\Delta t))/(2\Delta t)$, and the spatial derivatives are calculated with the Fourier pseudo-spectral method.
For the 3D turbulent scenario, we choose a forced isotropic turbulence simulation from the Johns Hopkins Turbulence Database (JHTDB) Databases (2019). It consists of a DNS simulation using a pseudo-spectral method with 10243 nodes in a periodic domain of size $(0,2\pi)^{3}$ over $T=10$ s, with 5028 timesteps. Energy is injected to maintain a constant total energy in modes such that their wave number magnitude is less than or equal to 2. We use a smaller derived dataset, taking 500 snapshots at a smaller resolution of $128^{3}$ using spectral downsampling. Finally, we shuffle all snapshots and use 400 and 100 samples for training and testing, respectively. The divergence of the flow field $\nabla\cdot\mathbf{u}$ is used as the metric for physics consistency. Figures 4(a) and 4(b) show samples of our training data.
@@ III.2 Physcial-consistency
We first target an unconditioned generation task to examine the ability of different methods to generate physically constrained samples. We consider standard diffusion models, Physics-Informed Diffusion Models Bastek et al. (2024) with dynamic rescale factor ($\text{PIDM}_{\text{dyn}}$), ConFIG Liu et al. (2024), and our $\text{ConFIG}_{\text{u}}$. Some comments about training hyperparameters and the network’s architecture can be found in Appendix B.
To calculate the physical residual, the time and spatial derivatives are evaluated in a discrete manner. Ideally, the physical residual of the training data, $\|\mathcal{R}(\omega^{*})\|_{2}$, should be 0. Still, we usually encounter a non-zero physical residual due to the discretisation errors. As no method can be expected to completely resolve this discretization error, it obscures the capabilities of the methods to improve the diffusion training. Hence, we employ a metric that compares the residual of the generated samples $\omega^{\prime}$ to the average of the test dataset samples:
@@
$\text{PIDM}_{\text{dyn}}$ Bastek et al. (2024) 0.99 $\pm$ 0.21 0.316 $\pm$ 0.005
ConFIG Liu et al. (2024) 1.24 $\pm$ 0.32 0.340 $\pm$ 0.009
@@
$\text{PIDM}_{\text{dyn}}$ Bastek et al. (2024) 1.08 $\pm$ 0.07 0.696 $\pm$ 0.042
ConFIG Liu et al. (2024) 1.01 $\pm$ 0.05 0.667 $\pm$ 0.067
@@ III.3 Reconstruction tasks
Next, we test the proposed masked diffusion method for reconstructing high-fidelity flow fields from sparse data points. Following the configuration from the previous study by Shu et al. Shu et al. (2023), we target reconstruction tasks with sparse observations of 5% and 1.5625% of the high fidelity data. An example of the sparse input can be seen in Figure 5.
Figure 5: From left to right: a sample of the dataset, and samples containing 5% and 1.5625% of the data, respectively.
@@
The first property is typically measured using the Root Mean Square Error (RMSE) between the predicted flow field and the ground truth. To satisfy physical coherence, the reconstructed field should follow the physical residual of the data, which can be evaluated using the relative residual loss $L_{\text{{res}}}$ introduced in Section III.2.
Evaluating the structure of a fluid flow improves our understanding of how closely the predicted sample aligns with the target high-fidelity field. As no ground truth is available in this probabilistic setting, we use a learned, perceptual similarity metric (LSiM) Kohl et al. (2020) to measure the similarity between our reconstruction and the ground truth. This metric allows us to assess whether the structural characteristics of the generated sample align with those of the target field.
@@
For the 2D scenario, we compute this error using the enstrophy spectrum instead of the energy spectrum, as used in other works Chen et al. (2024). This is because enstrophy cascades to small scales in 2D, providing a more sensitive and physically meaningful diagnostic of whether the simulation captures the correct multiscale turbulent dynamics.
@@
Besides comparing our method with that of Shu et al. Shu et al. (2023), we also compare our reconstruction result with a supervised baseline where the network directly predicts high-fidelity data from an interpolated sparse input. For a fair comparison, we use the same backbone architectures as the ones for the diffusion models. Moreover, we need to highlight that the baseline and other diffusion models for sparse data reconstruction require specific training for each new scenario (different % of data, for instance) to maximize their performance. In contrast, the masked diffusion method is general and only needs to be trained once, as an unconditional generator, using only the high-resolution training samples. Table 2 shows the results of the 2D 5% reconstruction task, where we can observe how our masked diffusion method outperforms previous methods on the RMSE, residual, structure similarity, and energy spectrum errors. We also provide the P-RMSE values, which measure the point-wise accuracy only at the sparse input data locations. These P-RMSE measurements highlight the improvements that masked diffusion provides at the locations constrained by input data.
Similar results can be observed when evaluating the 1.5625% task, which can be seen in Table 3. Our method consistently obtains the lowest errors in both tasks.
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.79 $\pm$ 0.09 0.79 $\pm$ 0.01 0.16 $\pm$ 0.07 0.29 $\pm$ 0.02 0.46 $\pm$ 0.12
Shu et al. Shu et al. (2023) 0.18 $\pm$ 0.03 0.14 $\pm$ 0.02 0.84 $\pm$ 0.07 0.10 $\pm$ 0.02 0.54 $\pm$ 0.08
Masked diffusion (ours) 0.13 $\pm$ 0.02 0.10 $\pm$ 0.01 0.02 $\pm$ 0.03 0.05 $\pm$ 0.01 0.26 $\pm$ 0.05
@@
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 1.06 $\pm$ 0.16 0.81 $\pm$ 0.02 0.21 $\pm$ 0.14 0.42 $\pm$ 0.06 0.68 $\pm$ 0.23
Shu et al. Shu et al. (2023) 0.37 $\pm$ 0.06 0.22 $\pm$ 0.03 0.22 $\pm$ 0.11 0.17 $\pm$ 0.03 0.52 $\pm$ 0.12
Masked diffusion (ours) 0.31 $\pm$ 0.06 0.15 $\pm$ 0.02 0.15 $\pm$ 0.06 0.11 $\pm$ 0.02 0.33 $\pm$ 0.07
@@
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.13 $\pm$ 0.01 0.22 $\pm$ 0.01 0.32 $\pm$ 0.02 0.05 $\pm$ 0.01 0.64 $\pm$ 0.02
Shu et al. Shu et al. (2023) 0.45 $\pm$ 0.03 0.68 $\pm$ 0.03 0.80 $\pm$ 0.06 0.64 $\pm$ 0.15 0.38 $\pm$ 0.07
Masked diffusion (ours) 0.17 $\pm$ 0.01 0.11 $\pm$ 0.01 0.61 $\pm$ 0.03 0.06 $\pm$ 0.01 0.05 $\pm$ 0.01
@@
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.16 $\pm$ 0.01 0.27 $\pm$ 0.01 0.36 $\pm$ 0.01 0.08 $\pm$ 0.01 0.71 $\pm$ 0.02
Shu et al. Shu et al. (2023) 0.44 $\pm$ 0.02 0.69 $\pm$ 0.04 0.79 $\pm$ 0.05 0.62 $\pm$ 0.11 0.36 $\pm$ 0.04
Masked diffusion (ours) 0.20 $\pm$ 0.01 0.18 $\pm$ 0.01 0.66 $\pm$ 0.03 0.09 $\pm$ 0.01 0.09 $\pm$ 0.01
@@
In Tables 4 and 5 we show the results for the 3D case of the same two reconstruction tasks. Our masked diffusion method outperforms Shu et al.’s method across all evaluation metrics. However, the supervised baseline achieves lower errors in RMSE, physical residual, and LSiM. These metrics are primarily influenced by the accuracy of the low-frequency components, which the supervised baseline captures more effectively. As shown in Figure 8, only the diffusion methods are capable of accurately reconstructing the high-frequency components. In contrast, the supervised baseline produces noticeably blurrier results despite achieving a lower RMSE. The spectrum in Figure 8 shows that the high wavenumber content of the solutions inferred by the supervised baseline clearly lacks energy. Another drawback of the supervised baseline (and other reconstruction methods like diffusion with conditioning on the low-resolution sample) is that it needs to be re-trained if the amount of available observations changes, while the method by Shu et al. Shu et al. (2023) and our masked diffusion only need to be trained once using only the high-resolution samples. Besides, as expected, our method achieves the lowest P-RMSE at the sparse data locations among the tested approaches. The masked diffusion method shown in these plots uses nearest interpolation to start the guidance in masked diffusion (our default case). However, one could use other, more advanced interpolation techniques for this purpose. We analyze an alternative in Appendix C.
@@
Figure 9: Example 2D Kolmogorov flow 5% data reconstruction.
Figure 10: Example 2D Kolmogorov flow 1.5625% data reconstruction.
Figure 11: Example of variability when reconstructing samples with 1.5625% of data. Mean and standard deviation were computed from 10 reconstructions.
@@
Figures 9-13 show samples of the reconstruction in 2D and 3D cases. It can be seen how masked diffusion better approximates the complicated structures of the turbulent flow fields, producing sharper images and yielding a closer approximation of the flow structures and statistics. In addition, since the diffusion model is inherently generative, it can produce diverse realizations for the same sparse input, as we show in Figure 11. The low deviation between samples highlights the high confidence of the model in the proposed reconstruction.
@@
As is evident from the spectrum errors in Tables 4 and 5, only the two considered diffusion models are capable of capturing small-scale features. E.g., our method features more than $6.4\times$ lower error than the supervised baseline. However, the two diffusion approaches behave very differently when considering different amounts of available data. In Figure 14, we plot the RMSE over the 3D reconstruction of 50 samples with Shu et al.’s method and our masked diffusion method (with nearest interpolation) together with direct linear interpolation as a baseline. Naturally, a reduced performance can be expected for very low amounts of input data. However, the two learned methods vary in their ability to handle the low-data regime. We observe that Shu et al.’s method remains mostly constant, with a consistently higher RMSE than our method. We see that if the percentage of available data is too low, our method achieves even lower RMSE errors than linear interpolation, in addition to producing high-frequency content. As a particularly challenging test case, table 6 shows the same metrics for a case with only 0.1% of sparse input data. Masked diffusion is the best at reconstructing the high-frequency components of the spectrum, as shown in Figure 15 for the challenging 0.1% data case. We include visualizations of several example reconstructions in Appendix D. Thus, the proposed method performs very well across the different regimes of data availability, for high percentage cases where linear interpolation and the supervised baseline still produce inaccurate high-wavenumber components, as well as highly sparse scenarios.
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Linear interpolation 0.33 $\pm$ 0.01 0.00 $\pm$ 0.01 1.75 $\pm$ 0.06 0.42 $\pm$ 0.04 0.75 $\pm$ 0.02
Supervised baseline 0.21 $\pm$ 0.01 0.36 $\pm$ 0.01 0.59 $\pm$ 0.01 0.17 $\pm$ 0.02 0.74 $\pm$ 0.02
Shu et al. Shu et al. (2023) 0.39 $\pm$ 0.04 0.69 $\pm$ 0.06 0.67 $\pm$ 0.03 0.46 $\pm$ 0.08 0.32 $\pm$ 0.02
Masked diffusion (ours) 0.27 $\pm$ 0.02 0.18 $\pm$ 0.01 0.74 $\pm$ 0.03 0.24 $\pm$ 0.08 0.23 $\pm$ 0.02
Table 6: Evaluation of the reconstruction task from 0.1% of the data for the 3D turbulent case.
@@
Figure 15: Energy spectrum comparison for a 0.1% 3D sparse data reconstruction.
We also study the distributional accuracy of the methods. For the 2D experiment, we measure the distribution of the vorticity magnitude of the reconstructed samples, as shown in Figure 16. Our method closely approximates the reference’s vorticity distribution on the 5% task, while Shu et al.’s method slightly underestimates the magnitude of the generated samples Shu et al. (2023). When evaluating the 1.5625% task, our method also underestimates the vorticity, but gives a better distribution than Shu et al.’s method.
@@
In Figure 17, we clearly observe that our masked diffusion is the closest to the ground truth, while the supervised baseline overestimates the zero component. Figures 18 show visualizations of the Q-criterion, highlighting how our proposed method is the only one that captures most of the smaller vortices in the fluid.
@@
Figure 18: Q-citerion in 3D for the 5% sparse data reconstruction. 3D plots showing Q isosurfaces (Q=80) with vorticity coloring.
When discussing efficiency, our method achieves these results using fewer reverse steps than Shu et al. Shu et al. (2023), making it a more computationally efficient model. For instance, for the more demanding 3D experiments in a NVIDIA L40S GPU, as shown in Appendix E, 100 reverse steps, i.e. NN evaluations, are enough to achieve maximum performance, compared to the T=1000 during training, or the 300 reverse steps used with Shu et al.’s method. This also results in shorter running times, where the supervised baseline requires around 0.04 s for a single sample, Shu et al.’s method takes 25 s, and our method takes 7.7 s. The only drawback of our method is a slight increase in memory demands. During masked diffusion, we need to store both the mask and the interpolated sparse data, resulting in 5.2 GB, compared with the 4.0 GB of Shu et al.’s method and the 1.4 GB of the single-step supervised baseline.
III.4 Ablation study
In this last section, we analyze the effect of different choices in our algorithm. First, the performance of diffusion models depends on the richness of our dataset Fukami and Taira (2024); Page (2025); Kadkhodaie et al. (2024); Gupta et al. (2025). We now examine the relationship between the reconstruction error and the number of training samples for our 3D masked diffusion reconstructions in Table 7. We present the metrics for models trained on the complete training dataset (100%) and reduced versions with 50% and 25% of the samples, all during the same number of iterations. While the physics residual remains unaffected, reduced data amounts result in increased LSiM distances and spectral errors. Training with the complete dataset provides better results and less variability.
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Masked diffusion (ours, default) 100% 0.20 $\pm$ 0.01 0.18 $\pm$ 0.01 0.66 $\pm$ 0.03 0.09 $\pm$ 0.01 0.09 $\pm$ 0.01
Masked diffusion (ours, default) 50% 0.22 $\pm$ 0.05 0.20 $\pm$ 0.08 0.66 $\pm$ 0.02 0.19 $\pm$ 0.16 0.21 $\pm$ 0.15
Masked diffusion (ours, default) 25% 0.22 $\pm$ 0.05 0.21 $\pm$ 0.06 0.71 $\pm$ 0.01 0.23 $\pm$ 0.17 0.24 $\pm$ 0.15
Table 7: Evaluation of the reconstruction task from 1.5625% of the data for the 3D turbulent case. We compare our masked diffusion reconstructions with models trained using 25%, 50% and 100% of the available training data.
Finally, we study the robustness of our method to noisy inputs. We evaluate the RMSE of our model on the 5% reconstruction task when Gaussian noise is added to the input. The model achieves its best accuracy when performing 100 neural network evaluations (as discussed in Appendix E) with noiseless input. However, in the presence of noise, we see a rapid decrease in accuracy when performing 100 reverse steps. We observe an inverse relationship between the ideal number of neural network evaluations and the amount of noise. In Figure 19, one can see how, when working with noisy inputs, the performance increases as the number of neural network evaluations decreases. Thus, the number of denoising steps can be interpreted as the strength of the guidance, which should be stronger when we are more certain about the precision of the input, and weaker when it is unreliable. Figure 19 shows that the number of reverse steps to perform a reconstruction should be chosen depending on the amount of noise that is present in the input.
Figure 19: RMSE on the 5% reconstruction task, where Gaussian noise $\sim\mathcal{N}(0,\sigma)$ is added to the input. The $\sigma$ is chosen to be a percentage of the dataset’s mean $\mu^{*}$.
Figure 20: Q-criterion distribution for the reconstructions of 3D TCF data with a 0.1% sparsity.
Figure 21: Energy spectrum for the reconstructions of 3D TCF data with 0.1% sparsity.
III.5 Summary of experiments and 3D outlook
Overall, our proposed training and sampling techniques demonstrate a clearly improved performance compared to existing approaches. These results highlight the ability of our masked diffusion framework to capture fine-grained details and preserve the structural integrity of the reconstructed fluid fields. To provide an outlook towards further 3D applications of our approach, we investigated the performance of our method on 3D turbulent channel flow (TCF) data at friction Reynolds number $Re_{\tau}=180$ Guastoni et al. (2020), obtaining results that align with the results reported above. We observed that the supervised baseline can achieve almost perfect reconstructions with 5% and 1.5% sparse data. However, the reconstruction performance degrades when less data is available (0.1%). In this case, our masked diffusion method performs much better, reconstructing most of the important turbulence scales, as shown in Figures 20 and 21.